# ADAM-SINDy: An Efficient Optimization Framework for Parameterized Nonlinear Dynamical System Identification

### v1

@@ 3.5 Pharmacokinetics Model — Time-dependent Reaction
This example investigates a dynamical system that incorporates a time-dependent function with a non-integer exponent in its governing equations, focusing on a pharmacokinetics model. Such models are commonly utilized to assess the rate of drug concentration across various compartments in the human body. Specifically, we consider a single-dose compartmental model 57 governed by the following system of equations:
@@
where $k_{g}$ denotes the time-dependent reaction rate 58, $k_{0}$ is the initial rate coefficient, and $\eta$ represents the fractal kinetics exponent, constrained within the range $10<\eta<$ 1. The ground truth data are generated using a drug absorption rate of $k_{b}=0.15$, $k_{0}=0.72$, and a fractal kinetics exponent of $\eta=-0.5$, evolved up to $t=10$ with the initial conditions $B_{0}=0$, $G_{0}=1$, and $U_{0}=0$ for a time step of $0.001$.
@@ 3.6 Wildfire Transport - Partial Differential Equation
The classical SINDy method was initially developed to discover systems of ODEs but has been extended to encompass PDEs through a process known as PDE-FIND 59. This extension is particularly pertinent for systems exhibiting dependencies in both spatial and temporal domains. In this section, we investigate wildfire dynamics modeled by PDEs, focusing on the identification of governing nonlinear parameters that elucidate the spatiotemporal evolution of temperature within the fire. These governing equations encapsulate critical processes, including heat transfer, combustion, and the influence of wind, thereby providing a comprehensive framework for understanding wildfire propagation across various fuel beds. The following wildfire combustion model 60 delineates this behavior, incorporating terms for advection, linear diffusion, reaction, and natural cooling:
@@
$\mathbf{\Theta}=\begin{bmatrix}1&\mathbf{X}&\nabla\mathbf{X}&\nabla^{2}\mathbf{X}&\nabla^{3}\mathbf{X}&\nabla^{4}\mathbf{X}&\exp\left(\frac{\mathbf{X}}{1+\mathcal{H}\mathbf{X}}\right)\end{bmatrix}\;,$ (15)
@@
3.7 Hyperparameter Robustness — Sparsity Knob $\lambda$
In this section, we investigate the optimization efficacy of the ADAM-SINDy framework in system identification by studying the effect of the sparsity-promoting parameters. To simplify the comparison with classical SINDy, $\mathbf{\Gamma}$ was fixed as the identity matrix, and the scalar $\lambda$ was made trainable. To study the sensitivity to this hyperparameter, an experiment using the SINDy-W method was conducted. By systematically adjusting the $\lambda$ parameter from smaller to larger values, the goal was to identify the threshold at which the model begins to diverge significantly from the ground truth. The objective was to estimate the minimum $\lambda$ ($\lambda_{\text{min}}$), below which the SINDy-W method generates an overly dense system of equations and the maximum $\lambda$ ($\lambda_{\text{max}}$), beyond which the model fails to capture meaningful dynamics. In ADAM-SINDy, $\lambda$ was initialized with these minimum and maximum threshold values and made trainable. All numerical experiments consistently demonstrated robust performance at the $\lambda_{\text{max}}$ initialized value, highlighting the ADAM-SINDy framework’s robustness. A notable example is the chemical kinetics problem, illustrated in Fig. 7 that demonstrates the robustness of ADAM-SINDy compared to SINDy-W, where the estimated values for $\lambda_{\text{max}}$ and $\lambda_{\text{min}}$ are $0.1$ and $0.001$, respectively. It could be seen that SIND-W can no longer produce the ground-truth equation when $\lambda_{\text{max}}$ was used, whereas ADAM-SINDy is more robust when initialized with $\lambda_{\text{max}}$. Furthermore, when $\lambda_{\text{min}}$ was used, both the SINDy-W and ADAM-SINDy methods failed (results not shown). This observation suggests that ADAM-SINDy demonstrates enhanced robustness in the selection of the sparsifying hyperparameter, particularly when initialized with larger values.
@@ 3.7 Hyperparameter Robustness — Sparsity Knob $\lambda$
Figure 7: The influence of the sparsity knob ($\lambda$) on chemical reaction kinetics results. (a) Loss plot showing the convergence of the ADAM-SINDy method for $\lambda_{\text{max}}$. A comparison of the governing equations is provided between the (b) ground truth and those identified by the (c) SINDy-W and (d) ADAM-SINDy approaches, for $\lambda_{\text{max}}=0.1$. The red-highlighted numerical values indicate the optimized parameters obtained.
@@ 4 Discussion
Identification of nonlinear dynamical systems is a prominent research domain that plays a pivotal role in elucidating the underlying physics of various problems. The literature highlights extensive investigations that employ the SINDy and symbolic regression frameworks, both of which can identify governing equations from observational data. However, it is essential to recognize that these well-established methods possess distinct advantages and disadvantages, particularly concerning computational optimization. In this context, the newly developed ADAM-SINDy framework emerges at the intersection of these established approaches, leveraging their computational optimization benefits. By employing the ADAM optimization algorithm, ADAM-SINDy facilitates the simultaneous optimization of parameters and coefficients associated with candidate nonlinear functions. This innovative approach enables more efficient and accurate parameter estimation without the necessity for prior knowledge of nonlinear characteristics, such as frequencies in trigonometric functions or bandwidths in exponential functions. A key factor in ADAM-SINDy’s success lies in its ability to exploit the computational graph inherent in the gradient descent algorithm implementation in the PyTorch environment. This structure facilitates efficient backpropagation of gradients, enabling dynamic interaction between candidate terms and nonlinear parameters during the optimization process. By allowing the model to adaptively refine both the nonlinear parameters and coefficients, ADAM-SINDy enhances the robustness and precision of system identification, particularly in complex dynamical systems where the classical SINDy method struggles with sparse, non-convex landscapes. Notably, the strategy behind the simultaneous optimization of all unknown parameters and coefficients closely parallels the use of adaptive activation functions 61 in neural network research, where the parameters of nonlinear adaptive activation functions, along with the weights and biases of linear layers, are optimized concurrently.
Interestingly, the ADAM-SINDy framework also facilitates interaction among the functions within the candidate library through the self-adaptation of nonlinear functions. This self-adaptive mechanism significantly enhances the model’s capacity to capture complex dynamical behaviors, thereby allowing it to adjust to the varying degrees of nonlinearity inherent in real-world systems. Such adaptability is particularly crucial when predefined candidates are lacking, as it optimizes parameters to compensate for absent functions. In contrast, the classical SINDy approach is constrained by static nonlinear parameters and a lack of interaction, increasing the risk of inaccurate system models when essential candidate functions are omitted. This adaptability parallels the advantages observed in symbolic regression techniques, further enriching the model’s applicability. Concurrently, ADAM-SINDy aims to identify a parsimonious and elegant mathematical model by incorporating essential computational methodologies, such as L1-norm loss functions and sequential thresholding, which are crucial for the success of the classical SINDy method. Furthermore, a distinct advantage of ADAM-SINDy lies in the optimization of sparsity-promoting hyperparameters such as $\mathbf{\Gamma}$. This optimization strategy reduces the need for extensive manual tuning, thereby enhancing the model’s robustness. This trainable $\mathbf{\Gamma}$ further advances the Adaptive LASSO method, by enabling the dynamic imposition of candidate-specific penalties during the optimization process, leading to a “fully” adaptive LASSO approach. Such adaptive weighting/scaling flexibility is particularly advantageous in applications involving high-dimensional datasets, as it reduces the risk of overfitting while simultaneously enhancing the interpretability of the model. While ADAM-SINDy offers both batch-wise and full-batch optimization, it has been observed that for simpler dynamical systems, a small randomly sampled batch of data is often sufficient to capture the necessary temporal information for accurately estimating governing parameters. However, for more complex systems, a small batch may not encompass all required information, necessitating a larger batch size that ideally approaches full-batch size. Consequently, full-batch optimization is preferred and utilized in the present work, leading to reduced uncertainty in the converged optimized parameters within the ADAM-SINDy methodology.
@@
Ultimately, the integrated simultaneous adaptive optimization of the parameters of nonlinear candidate functions, coefficients of the candidate matrix, and hyperparameters based on training data positions ADAM-SINDy as a powerful and effective tool for modeling a diverse range of dynamical behaviors across various applications. The proposed ADAM-SINDy methodology is thoroughly investigated across a wide spectrum of dynamical systems, including benchmark coupled ODEs such as oscillators, fluid flows, reaction kinetics, and wildfire dynamics PDE. These comprehensive investigations demonstrated its efficacy in accurately computing nonlinear features such as multiple Fourier frequencies, non-integer exponents in polynomials, and exponential exponents. Furthermore, the investigation also explored the inherent limitations of the classical SINDy method, illustrating its consistent failure to identify the exact model when the precise parameters are absent from the candidate matrix. This comprehensive investigation highlights the considerable potential of the ADAM-SINDy optimization framework, thereby extending its applicability to the modeling of intricate and diverse dynamical systems across various contexts with limited prior knowledge.
It is worth noting that recent advancements in symbolic regression via neural networks have focused on identifying dynamical systems 55. Boddupalli et al. 55 applied this approach to chemical reaction kinetics, the same system considered in our work (Sec. 3.4), and their results produced overly complex, “bloated” equations characterized by cubic polynomials nonlinearly combined with exponential functions 55. While the inclusion of stacked exponential terms alongside polynomials was necessary for convergence, it significantly increased model complexity, leading to inaccuracies in predicting the true trajectory and governing dynamics 55. In contrast to the considered growth rates of unity for reactant concentration and temperature 55, 56, we applied ADAM-SINDy to the same chemical reaction system (see Fig. 4), tackling a more computationally challenging scenario involving nonlinear parameter estimation with non-integer rational exponents of $1.015$ and $1.025$ for reactant concentration and temperature, respectively. Remarkably, ADAM-SINDy accurately identified the true dynamical system, capturing both slow and fast transients without requiring additional treatments to ensure regression convergence.
@@
In the context of the ABC flow example, efforts were made to leverage the computational graph of the gradient descent algorithm by incorporating a divergence-free constraint into the loss equation. This approach demonstrated a noteworthy reduction in the total number of epochs required for optimization—approximately 20 epochs—compared to the typical requirement of around $6\times 10^{4}$ epochs. This observation highlights the potential for further enhancement of the method in future work. Thus, a promising direction for future research involves exploring differential operator-based sparsity controllers ($\mathbf{\Gamma}$). Such advancements could facilitate the development of a more parsimonious and interpretable dynamical system model that aligns with underlying physical laws, thereby significantly improving the computational efficacy of the ADAM-SINDy framework.
@@ Acknowledgments
This research was supported by the National Science Foundation under Award No. 2247173 and 2330212. The authors would like to express their gratitude to Dr. Jacob Hochhalter and Dr. Mike Kirby for their valuable discussions about this work.

### v2

@@ 3.5 Pharmacokinetics Model — Time-dependent Reaction
This example investigates a dynamical system that incorporates a time-dependent function with a non-integer exponent in its governing equations, focusing on a pharmacokinetics model. Such models are commonly utilized to assess the rate of drug concentration across various compartments in the human body. Specifically, we consider a single-dose compartmental model [62] governed by the following system of equations:
@@
where $k_{g}$ denotes the time-dependent reaction rate [63], $k_{0}$ is the initial rate coefficient, and $\eta$ represents the fractal kinetics exponent, constrained within the range $10<\eta<$ 1. The ground truth data are generated using a drug absorption rate of $k_{b}=0.15$, $k_{0}=0.72$, and a fractal kinetics exponent of $\eta=-0.5$, evolved up to $t=10$ with the initial conditions $B_{0}=0$, $G_{0}=1$, and $U_{0}=0$ for a time step of $0.001$.
@@ 3.6 Wildfire Transport - Partial Differential Equation
The classical SINDy method was initially developed to discover systems of ODEs but has been extended to encompass PDEs through a process known as PDE-FIND [64]. This extension is particularly pertinent for systems exhibiting dependencies in both spatial and temporal domains. In this section, we investigate wildfire dynamics modeled by PDEs, focusing on the identification of governing nonlinear parameters that elucidate the spatiotemporal evolution of temperature within the fire. These governing equations encapsulate critical processes, including heat transfer, combustion, and the influence of wind, thereby providing a comprehensive framework for understanding wildfire propagation across various fuel beds. The following non-dimensional wildfire combustion model [65] delineates this behavior, incorporating terms for advection, linear diffusion, reaction, and natural cooling:
@@
$\bm{\Theta}=\begin{bmatrix}1&\mathbf{X}&\nabla\mathbf{X}&\nabla^{2}\mathbf{X}&\nabla^{3}\mathbf{X}&\nabla^{4}\mathbf{X}&\exp\left(\frac{\mathbf{X}}{1+\mathcal{H}\mathbf{X}}\right)\end{bmatrix}\;,$ (17)
@@
3.7 Hyperparameter Robustness — Sparsity Knobs
In this section, we investigate the optimization efficacy of the ADAM-SINDy framework in system identification, focusing on its sensitivity to the sparsity-promoting hyperparameters, $\bm{\Gamma}$ and $\lambda$. The performance of ADAM-SINDy is compared with the ideal case, the SINDy-W method. First, the SINDy-W method is employed to systematically vary the $\lambda$ parameter across one order of magnitude, ranging from smaller to larger values, in order to determine the bounds beyond which the model diverges significantly from the ground truth. Specifically, we identify the minimum $\lambda$, below which the SINDy-W method generates an overly dense system of equations, and the maximum $\lambda$, beyond which the model fails to capture meaningful dynamics. The computed bound values for the SINDy-W method are presented in Table 2.
In the context of the ADAM-SINDy method, two scenarios are considered: $(i)$ $\bm{\Gamma}$ is fixed to the value of 1, while $\lambda$ is trainable, and $(ii)$ $\bm{\Gamma}$ is trainable, with its initialization varied by varying the standard deviation ($\bm{\Gamma}_{\sigma}$) of a Gaussian distribution while keeping its mean ($\bm{\Gamma}_{\mu}$) at zero, and $\lambda$ is fixed at a value of 1. Experiments are conducted to compute the bounds for both scenarios across all the numerical examples considered in the present work, and the corresponding bounds are tabulated in Table 2. Generally, the ADAM-SINDy framework provides much wider bounds for the sparsity knobs, often at least one order magnitude higher, especially in cases with higher initialization values, indicating that ADAM-SINDy is more robust than the classical SINDy method. Comparing the trainable $\lambda$ and trainable $\bm{\Gamma}$ cases in ADAM-SINDy, it is observed that in each example one of the two approaches provides a wider bound for the sparsity knob, indicating no clear advnatage. However, in the former scenario (trainable $\lambda$), ADAM-SINDy identifies the correct candidate terms but the coefficients and nonlinear parameters are not always machine precision. In the latter case (trainable $\bm{\Gamma}$), the ADAM-SINDy method identifies the exact correct terms and nonlinear parameter values with machine precision.
Our experiments demonstrate that in general the latter optimization strategy, which imposes candidate-wise sparsity, is recommended for nonlinear system identification.
A notable example is the chemical kinetics problem, illustrated in Fig. 7, which represents the former scenario of trainable $\lambda$. It is observed that the SINDy-W method fails to produce the ground-truth equation, whereas ADAM-SINDy remains robust, producing the correct terms with slight deviations in the parameter values when initialized with a higher $\lambda$ value of 0.1. Furthermore, results for the latter scenario with trainable $\bm{\Gamma}$ are not shown here, as the identified terms and parameter values are observed to match the ground truth exactly, as demonstrated in Fig. 4.
@@ 3.7 Hyperparameter Robustness — Sparsity Knobs
Figure 7: The influence of the sparsity knob ($\lambda$) on chemical reaction kinetics results. (a) Loss plot showing the convergence of the ADAM-SINDy method for $\lambda=0.1$. A comparison of the governing equations is provided between the (b) ground truth and those identified by the (c) SINDy-W and (d) ADAM-SINDy methods, for $\lambda=0.1$. The red-highlighted numerical values indicate the optimized parameters obtained.
Experiment SINDy-W
ADAM-SINDy
$\lambda$
$\bm{\Gamma}$
Harmonic oscillator
$10^{-3}$ $<$ $\lambda$ $<$ $10^{-1}$
$10^{-3}$ $<$ $\lambda$ $<$ 10
$10^{-8}$ $<$ $\bm{\Gamma}_{\sigma}$ $<$ 10
Van der pol oscillator
$10^{-3}$ $<$ $\lambda$ $<$ $10^{-2}$
$10^{-5}$ $<$ $\lambda$ $<$ 10
$10^{-5}$ $<$ $\bm{\Gamma}_{\sigma}$ $<$ 10
ABC flow
$10^{-1}$
$10^{-1}$
$10^{-2}$ $<$ $\bm{\Gamma}_{\sigma}$ $<$ 10
Chemical reaction
$10^{-3}$ $<$ $\lambda$ $<$ $10^{-1}$
$10^{-12}$ $<$ $\lambda$ $<$ 10
$10^{-9}$ $<$ $\bm{\Gamma}_{\sigma}$ $<$ 10
Pharmacokinetics
$10^{-2}$ $<$ $\lambda$ $<$ $10^{-1}$
$10^{-3}$ $<$ $\lambda$ $<$ 10
1 $<$ $\bm{\Gamma}_{\sigma}$ $<$ 10
Wildfire PDE
$10^{-6}$ $<$ $\lambda$ $<$ 1
$10^{-6}$ $<$ $\bm{\Gamma}_{\sigma}$ $<$ 1
Table 2: Bounds on the sparsity-promoting hyperparameters for the SINDy-W and ADAM-SINDy methodologies, illustrating the conditions under which each method accurately identifies the dynamical system. In the ADAM-SINDy method, $\lambda$ is trainable while $\bm{\Gamma}$ is fixed at 1 in the first column. It is noted that no $\lambda$ values were found to work for the wildfire PDE example. In the second column, $\bm{\Gamma}$ is trainable while $\lambda$ is fixed at 1 and $\bm{\Gamma}$ is initialized using a Gaussian distribution where its mean $\bm{\Gamma}_{\mu}$ is fixed at zero and its standard deviation $\bm{\Gamma}_{\sigma}$ is varied as shown.
@@ 4 Discussion
Identification of nonlinear dynamical systems is a prominent research domain that plays a pivotal role in elucidating the underlying physics of various problems. The literature highlights extensive investigations that employ the SINDy and symbolic regression frameworks, both of which can identify governing equations from observational data. However, it is essential to recognize that these well-established methods possess distinct advantages and disadvantages, particularly concerning computational optimization. In this context, the newly developed ADAM-SINDy framework emerges at the intersection of these established approaches, leveraging their computational optimization benefits. By employing the ADAM optimization algorithm, ADAM-SINDy facilitates the simultaneous optimization of parameters and coefficients associated with candidate nonlinear functions. This innovative approach enables more efficient and accurate parameter estimation without the necessity for prior knowledge of nonlinear characteristics, such as frequencies in trigonometric functions or bandwidths in exponential functions. A key factor in ADAM-SINDy’s success lies in its ability to exploit the computational graph inherent in the gradient descent algorithm implementation in the PyTorch environment. This structure facilitates efficient backpropagation of gradients, enabling dynamic interaction between candidate terms and nonlinear parameters during the optimization process. By allowing the model to adaptively refine both the nonlinear parameters and coefficients, ADAM-SINDy enhances the robustness and precision of system identification, particularly in complex dynamical systems where the classical SINDy method struggles with sparse, non-convex landscapes. Notably, the strategy behind the simultaneous optimization of all unknown parameters and coefficients closely parallels the use of adaptive activation functions [66] in neural network research, where the parameters of nonlinear adaptive activation functions, along with the weights and biases of linear layers, are optimized concurrently.
Interestingly, the ADAM-SINDy framework also facilitates interaction among the functions within the candidate library through the self-adaptation of nonlinear functions. This self-adaptive mechanism significantly enhances the model’s capacity to capture complex dynamical behaviors, thereby allowing it to adjust to the varying degrees of nonlinearity inherent in real-world systems. Such adaptability is particularly crucial when predefined candidates are lacking, as it optimizes parameters to compensate for absent functions. In contrast, the classical SINDy approach is constrained by static nonlinear parameters and a lack of interaction, increasing the risk of inaccurate system models when essential candidate functions are omitted. This adaptability parallels the advantages observed in symbolic regression techniques, further enriching the model’s applicability. However, the occurrence of self-adaptation may transform the candidate matrix prior to optimization into an “ill-conditioned” matrix following optimization. Based on our tests, ill-conditioned cases do not pose a problem for ADAM-SINDy and simply result in two identical terms (in different forms) that need to be manually added for simplification of the final equation (see Fig. 2).
Concurrently, ADAM-SINDy aims to identify a parsimonious mathematical model while optimizing trainable sparsity-promoting hyperparameters, such as $\bm{\Gamma}$. A similar adaptive weighting strategy can be found in other research works, such as PINN [57] and Transformers [67]. From the sensitivity analysis in Sec. 3.7, it is evident that both trainable hyperparameters can be employed for ADAM-SINDy optimization. However, prioritizing the trainable $\bm{\Gamma}$ is recommended, as it reliably produces the correct terms with machine precision, whereas $\lambda$ may identify the correct terms but with slight deviations in their numerical values. While ADAM-SINDy offers both batch-wise and full-batch optimization, it has been observed that for simpler dynamical systems, a small randomly sampled batch of data is often sufficient to capture the necessary temporal information for accurately estimating governing parameters. However, for more complex systems, a small batch may not encompass all required information, necessitating a larger batch size that ideally approaches full-batch size. Consequently, full-batch optimization is preferred and utilized in the present work, leading to reduced uncertainty in the converged optimized parameters within the ADAM-SINDy methodology.
@@
Ultimately, the integrated simultaneous adaptive optimization of the nonlinear parameters, coefficients matrix, and hyperparameters based on training data positions ADAM-SINDy as a powerful and effective tool for modeling a diverse range of dynamical behaviors across various applications. The proposed ADAM-SINDy methodology is thoroughly investigated across a wide spectrum of dynamical systems, including benchmark coupled ODEs such as oscillators, fluid flows, reaction kinetics, and wildfire dynamics PDE. These comprehensive investigations demonstrated its efficacy in accurately computing nonlinear features such as multiple Fourier frequencies, non-integer exponents in polynomials, and exponential exponents. Furthermore, the investigation also explored the inherent limitations of the classical SINDy method, illustrating its consistent failure to identify the exact model when the precise parameters are absent from the candidate matrix. This comprehensive investigation highlights the considerable potential of the ADAM-SINDy optimization framework, thereby extending its applicability to the modeling of intricate and diverse dynamical systems across various contexts with limited prior knowledge.
It is worth noting that symbolic regression via neural networks has been employed to identify dynamical systems [60]. Boddupalli et al. [60] applied this approach to chemical reaction kinetics, the same system considered in this work (Sec. 3.4), resulting in an overly complex, “bloated” mathematical model when explicit stacking within the neural network is absent. However, when exponential terms are explicitly stacked with polynomials, the identified model closely approximates the ground truth equation but exhibits slight deviations in the coefficient values. Remarkably, ADAM-SINDy accurately identified the true dynamical system, capturing the dynamics without requiring additional treatments to ensure regression convergence.
@@
In the context of the ABC flow example, efforts were made to leverage the computational graph of the gradient descent algorithm by incorporating a divergence-free constraint into the loss equation. This approach demonstrated a noteworthy reduction in the total number of epochs required for optimization—approximately 40 epochs—compared to the typical requirement of around $6\times 10^{4}$ epochs. This observation highlights the potential for further enhancement of the method in future work. Thus, a promising direction for future research involves exploring differential equation-based sparsity controllers. Such advancements could facilitate the development of a more parsimonious and interpretable dynamical system model that aligns with underlying physical laws, thereby significantly improving the computational efficacy of the ADAM-SINDy framework.
@@ Acknowledgments
This research was supported by the National Science Foundation under Award No. 2247173 and 2330212. The authors express their gratitude to Dr. Jacob Hochhalter, Dr. Sarang Joshi, and Dr. Mike Kirby for their valuable discussions about this work.